Determine the following integrals:

1. "\smallint" [x-2/x^2][X+2/x^2]DX

2. "\smallint" e^5x[e^2x/7+3/e^3x]DX

3."\smallint" 1/(4-√3x)^3 DX

4. "\smallint" π/40 (tan X)^3 (sec X )^3 dx





1
Expert's answer
2021-09-05T17:56:27-0400

"1. \int [x-\frac{2}{x^2}][x+\frac{2}{x^2}]dx\\
=\int [x^2-\frac{4}{x^4}]dx\\
= \frac{x^3}{3}+\frac{4}{3x^3}+c\\


2. \int e^{5x}[\frac{e^{2x}}{7}+\frac{3}{e^{3x}}]dx\\
=\int [\frac{e^{7x}}{7}+3{e^{2x}}]dx\\
=\frac{e^{7x}}{49}+\frac{3e^{2x}}{2}+c\\



3.\int \frac{1}{(4-\sqrt{3x})^3}dx\\
\text{put}\\
4-\sqrt{3x}=t\\
=>\frac{-3}{2\sqrt{3x}}dx=dt\\
\implies dx=\frac{-2}{3}(4-t)dt\\
\int \frac{1}{(4-\sqrt{3x})^3}dx\\
=\frac{-2}{3} \int \frac{4-t}{t^3}dt\\
=\frac{-2}{3} \int (4t^{-3}-t^{-2})dt\\
=\frac{-2}{3}(\frac{-2}{t^{2}}+\frac{1}{3t^3})+c\\
=\frac{-2}{3}(\frac{-2}{(4-\sqrt{3x})
^{2}}+\frac{1}{3(4-\sqrt{3x})^3})+c\\

4. \int^{π/4}_{0 } (tan x)^3 (sec x )^3dx\\ 
\text{put},\\
secx=t\\
=>tanx \space secx \space
 dx=dt\\
\text{and using }tan^2x=1-sec^2x,
\text{we get}\\
\int^{π/4}_{0 } (tan x)^3 (sec x )^3dx\\
=\int^{π/4}_{0 } (secx)^2(tan x )^2tanx \space secx\space dx\\
=\int^{π/4}_{0} (sec x )^2(1-sec^2x)tanx \space secx\space dx\\
 =\int^{\sqrt2}_1 t^2(1-t^2)dt\\
=\int^{\sqrt2}_1 (t^2-t^4)dt\\
=[\frac{t^3}{3}-\frac{t^5}{5}]^{\sqrt2}_1\\
=-0.32"